# Graded Projection Recursion (GPR): A Framework for Controlling Bit-Complexity of Algebraic Packing

### v2

@@ Abstract
We present Graded Projection Recursion (GPR), a framework for converting certain blocked recursive algorithms into model-honest (i.e., reflects full bit complexity) near-quadratic procedures under bounded intermediate budgets. Part I gives a proof-complete integral specification of a three-band packing identity and a two-round middle-band extractor, and shows how per-node centering and sqrt-free dyadic $\ell_{2}$ normalization (recursive invariant amplification) gives a sufficient packing base $\beta(\sigma)$ that grows linearly with the scaling depth $\sigma$ (i.e., logarithmic bit-growth) in exact arithmetic.
@@ 1 Executive summary: why GPR avoids degree-mixing and bit-growth obstructions
    Depth-uniform boundedness. If graded projection gives a depth-uniform bound on the target/tail bands (Part IV master condition or BWBM), then a choice of extractor parameters dependent only on depth (e.g., $\beta\propto 2^{\sigma}$) suffices.
@@ 3.1 Dyadic lattice and scaled coefficient lattice
Definition 3.2 (Scaled coefficient lattice).
Recursive invariant amplification introduces dyadic scalings at each node. To maintain exactness at a recursion node with cumulative scaling exponent $\sigma\in\mathbb{Z}_{\geq 0}$, we define the local scaled coefficient lattice:
@@
The rounding operator relative to this lattice is defined via the lifted integers:
@@
If $z\in\mathcal{L}_{\mathrm{coeff}}(\sigma)$, then $\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil=z$ exactly.
@@ 5.1 Two-round sieve on scaled lattices
$\left\lfloor Mz/\beta\right\rceil~=~MU\quad\text{and}\quad\left\lfloor Mz\right\rceil~=~\beta(MU)+MT.$ (15)
@@ 6.1 Dyadic $\ell_{2}$ normalization without square roots
Definition 6.1 (Lifted squared norm).
@@
$S(x)~:=~\sum_{r=1}^{m}\left(Dx_{r}\right)^{2}\in\mathbb{Z}_{\geq 0}.$ (34)
Equivalently, if $x_{r}=(w_{r}/D)2^{e_{r}}$ with $w_{r}\in\mathbb{Z}$, then $Dx_{r}=w_{r}2^{e_{r}}$ is dyadic and $S(x)$ is the exact squared $\ell_{2}$ norm of the lifted vector $Dx$.
@@
Let $x\in\mathcal{L}_{D}^{m}$ and define $S(x)\in\mathbb{Z}_{\geq 0}$ and $k(x)\in\mathbb{Z}_{\geq 0}$ as in Definitions 6.16.2, then:
@@
    (No square roots needed.) If $S(x)>0$ and $\ell:=\lfloor\log_{2}S(x)\rfloor$ (bitlength minus one), let $k_{0}:=\lfloor\ell/2\rfloor$, then $k(x)$ is either $k_{0}$ or $k_{0}+1$, i.e.:
    $k(x)~=~\begin{cases}k_{0},&\text{if}~S(x)\leq 2^{2k_{0}},\\ k_{0}+1,&\text{otherwise.}\end{cases}$ (37)
    Equivalently, $k(x)=\left\lceil\frac{1}{2}\log_{2}S(x)\right\rceil$ with the convention $k(0)=0$.
@@ 7 Sufficient-base theorem
Fix $\delta\in(0,\frac{1}{2}]$. Assume that at a recursion node with cumulative scaling $\sigma$, the normalized operands satisfy the dimension-free dot-product bounds $|T_{ij}|\leq 1$ and $|L_{ij}|\leq 1$.
@@
Remark 7.2 (Why a growing base is not “cheating”).
Condition (48) implies that $\beta$ must grow roughly in proportion to $2^{\sigma}$. Because $\sigma$ represents the cumulative normalization depth, $\sigma\approx\log n$. Consequently, the required bit-width of $\beta$ grows logarithmically ($O(\log n)$). This is not a violation of the model; rather, it is the precise origin of the digit-band staging factor $K(n)$ (Theorem 9.3). The algorithm exchanges the uncontrolled magnitude growth of classical methods for a strictly controlled, logarithmic growth in precision (staging depth), giving the soft-quadratic complexity bound.
@@ 8.1 Centering correction identity
Let $A,B\in\mathbb{R}^{n\times n}$ and let $\alpha,\beta\in\mathbb{R}^{n}$, and define the centered matrices:
@@
$AB~=~A_{0}B_{0}+\mathbf{1}(\alpha^{\top}B_{0})+(A_{0}\beta)\mathbf{1}^{\top}+(\alpha^{\top}\beta)\mathbf{1}\mathbf{1}^{\top}.$ (51)
@@
From the definitions, $A=A_{0}+\mathbf{1}\alpha^{\top}$ and $B=B_{0}+\beta\mathbf{1}^{\top}$. Multiply and expand:
$AB$ $=$ $(A_{0}+\mathbf{1}\alpha^{\top})(B_{0}+\beta\mathbf{1}^{\top})$ (52)
$=$ $A_{0}B_{0}+A_{0}(\beta\mathbf{1}^{\top})+(\mathbf{1}\alpha^{\top})B_{0}+(\mathbf{1}\alpha^{\top})(\beta\mathbf{1}^{\top}).$ (53)
Now simplify each term:
$A_{0}(\beta\mathbf{1}^{\top})$ $=$ $(A_{0}\beta)\mathbf{1}^{\top},$ (54)
@@
and:
$(\mathbf{1}\alpha^{\top})(\beta\mathbf{1}^{\top})$ $=$ $\mathbf{1}(\alpha^{\top}\beta)\mathbf{1}^{\top}$ (56)
$=$ $(\alpha^{\top}\beta)\mathbf{1}\mathbf{1}^{\top}.$ (57)
@@
We refer to this as the row(A)-col(B) policy. It implies the entrywise bound $|(A^{\prime}B^{\prime})_{ij}|\leq 1$ by Cauchy–Schwarz, which is the invariant used to justify a constant packing base $\beta_{\star}$.
@@
Centering enforces mean constraints. Normalization enforces norm constraints. These operations generally do not commute, and we cannot in general enforce zero-mean rows and columns and unit row and column norms simultaneously. Throughout, we use the row(A)–col(B) policy (Definition 8.2) specifically to guarantee the entrywise bound $|(A^{\prime}B^{\prime})_{ij}|\leq 1$ needed for a constant $\beta_{\star}$.
@@
This choice is deliberate because it implies the entrywise bound $|(A^{\prime}B^{\prime})_{ij}|\leq 1$, which is the invariant needed to justify a constant packing base $\beta_{\star}$.
@@ 8.2 Implementation contract (integral kernel)
    No square roots: Dyadic $\ell_{2}$ normalization uses $S(x)=\sum(Dx_{r})^{2}$ and $k(x)$ via bitlength comparisons (Proposition 6.3). Scaling by $2^{\pm k}$ is an exponent-offset update.
@@
    Sufficient base: The packing base $\beta$ is chosen to satisfy Theorem 7.1 given the node’s scaling $\sigma$ (e.g., via the allocated staging budget).
@@
    Shifts do not touch mantissas: Multiplication by $2^{\pm t}$ and by $\beta_{\star}^{\pm 1}$ must be implemented as exponent updates, not as integer multiplication on mantissas.
@@ 8.3 Kernel pseudocode
(1) Center. Compute the exact column-mean vector $\alpha\in\mathcal{L}_{D}^{n}$ of $A$ and exact row-mean vector $\beta\in\mathcal{L}_{D}^{n}$ of $B$ using dyadic division:
@@
$\beta_{i}$ $:=$ $\frac{1}{n}\sum_{j=1}^{n}B_{ij}.$
Set $A_{0}:=A-\mathbf{1}\alpha^{\top}$ and $B_{0}:=B-\beta\mathbf{1}^{\top}$.
@@
(3) Pack, recurse, and extract quadrants. Partition $A^{\prime},B^{\prime}$ into $2\times 2$ blocks of size $(n/2)\times(n/2)$. Choose base $\beta$ satisfying Theorem 7.1 given current scaling $\sigma$. foreach $(i,j)\in\{1,2\}^{2}$ do
@@
$C~:=~A_{0}B_{0}+\mathbf{1}(\alpha^{\top}B_{0})+(A_{0}\beta)\mathbf{1}^{\top}+(\alpha^{\top}\beta)\mathbf{1}\mathbf{1}^{\top}.$

### v3

@@ Abstract
We present Graded Projection Recursion (GPR), a framework for converting certain blocked recursive algorithms into model-honest (i.e., reflects full bit complexity) near-quadratic procedures under bounded intermediate budgets. Part I gives a proof-complete integral specification of a three-band packing identity and a two-round middle-band extractor, and shows how per-node centering and sqrt-free dyadic $\ell_{2}$ normalization (recursive invariant amplification) gives a sufficient packing base $\beta(\sigma)$ whose required bit-width grows linearly with the node-local lattice exponent $\sigma$ (so only logarithmic bit-growth in $n$ in the worst case) in exact arithmetic.
@@ 1 Executive summary: why GPR avoids degree-mixing and bit-growth obstructions
    Depth-uniform boundedness. If graded projection gives a depth-uniform bound on the target/tail bands (Part IV master condition or BWBM), then a choice of extractor parameters dependent only on the node-local lattice exponent $\sigma$ (e.g., $\beta=\beta(\sigma)$) suffices.
@@ 3.1 Dyadic lattice and scaled coefficient lattice
Definition 3.2 (Scaled coefficient lattice and node-local lattice exponent).
Fix $D\in\mathbb{N}$ and recall $\mathcal{L}_{D}=D^{-1}\mathbb{Z}[1/2]$ (Definition 3.1). At each recursion node (dimension $m$), after centering and dyadic $\ell_{2}$ normalization have produced normalized operands $A^{\prime},B^{\prime}\in\mathcal{L}_{D}^{m\times m}$, we define the node-local coefficient-lattice exponent $\sigma\in\mathbb{Z}_{\geq 0}$ as the smallest integer such that all dot-product terms $A^{\prime}_{ir}B^{\prime}_{rj}$ lie in $D^{-2}2^{-\sigma}\mathbb{Z}$.
Write each stored scalar in mantissa/exponent-offset form $x=(w/D)2^{e}$ (Appendix B.1) and denote its exponent offset by $\exp(x):=e$, with the convention $\exp(0):=+\infty$. Define for each inner index $r\in\{1,\dots,m\}$:
$a_{r}$ $:=$ $\min_{1\leq i\leq m}\exp(A^{\prime}_{ir}),$ (2)
$b_{r}$ $:=$ $\min_{1\leq j\leq m}\exp(B^{\prime}_{rj}),$ (3)
and let:
$\mathcal{R}~:=~\{r\in\{1,\dots,m\}:a_{r}<\infty~\text{and}~b_{r}<\infty\}.$ (4)
If $\mathcal{R}=\emptyset$ (i.e., for every $r$ either $A^{\prime}_{:,r}\equiv 0$ or $B^{\prime}_{r,:}\equiv 0$, so all dot-product terms vanish), then set $\sigma:=0$. Otherwise define:
$g_{\min}~:=~\min_{r\in\mathcal{R}}(a_{r}+b_{r}),\\ ~~~\sigma:=\max\{0,\,-g_{\min}\}.$ (5)
The scaled coefficient lattice at this node is then:
@@
With this choice of $\sigma$, every band coefficient $U,T$ in the three-band identity (Proposition 5.7) for products of normalized blocks lies in $\mathcal{L}_{\mathrm{coeff}}(\sigma)$, so the scaled rounding operator is well-defined:
@@
The scaled sieve extractor is:
$\mathcal{S}_{\beta,\sigma}(z)~:=~\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil-\beta\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z/\beta\right\rceil$ (8)
(cf., Definition 5.1 and Appendix B.6).
@@ 5.1 Two-round sieve on scaled lattices
$\left\lfloor Mz/\beta\right\rceil~=~MU~\text{and}~\left\lfloor Mz\right\rceil~=~\beta(MU)+MT.$ (20)
@@ 6.1 Dyadic $\ell_{2}$ normalization without square roots
Definition 6.1 (Lifted squared norm (dyadic form)).
@@
$S(x)~:=~\sum_{r=1}^{m}\left(Dx_{r}\right)^{2}\in\mathbb{Z}[1/2]_{\geq 0}.$ (39)
Equivalently, if $x_{r}=(w_{r}/D)2^{e_{r}}$ with $w_{r}\in\mathbb{Z}$ and $e_{r}\in\mathbb{Z}$, then $Dx_{r}=w_{r}2^{e_{r}}$ and $S(x)=\sum_{r}w_{r}^{2}2^{2e_{r}}$.
For implementation it is convenient to form an integer lift. Let $t:=\max\{0,-\min_{r}e_{r}\}\in\mathbb{Z}_{\geq 0}$ and define
$\widetilde{S}(x)~:=~\sum_{r=1}^{m}\bigl(w_{r}2^{e_{r}+t}\bigr)^{2}\in\mathbb{Z}_{\geq 0},$ (40)
then $S(x)=2^{-2t}\widetilde{S}(x)$ exactly.
@@
Let $x\in\mathcal{L}_{D}^{m}$ and define its lifted squared norm $S(x)\in\mathbb{Z}[1/2]_{\geq 0}$ and integer lift $\widetilde{S}(x)\in\mathbb{Z}_{\geq 0}$ as in Definition 6.1. Let $k(x)\in\mathbb{Z}_{\geq 0}$ be the dyadic scaling exponent from Definition 6.2, then:
@@
    (No square roots needed (via an integer lift).) If $\widetilde{S}(x)>0$ and $\ell:=\lfloor\log_{2}\widetilde{S}(x)\rfloor$ (bitlength minus one), let $K_{0}:=\lfloor\ell/2\rfloor$ and set:
    $K(x)~=~\begin{cases}K_{0},&\text{if}~\widetilde{S}(x)\leq 2^{2K_{0}},\\ K_{0}+1,&\text{otherwise,}\end{cases}~~k(x)~=~\max\{0,\,K(x)-t\},$ (43)
    where $t:=\max\{0,-\min_{r}e_{r}\}$ is the shift from Definition 6.1 satisfying $S(x)=2^{-2t}\widetilde{S}(x)$. Equivalently, $k(x)=\max\{0,\,\lceil\frac{1}{2}\log_{2}\widetilde{S}(x)\rceil-t\}$, with $k(0)=0$.
@@ 7 Sufficient-base theorem
Fix $\delta\in(0,\frac{1}{2}]$. Assume that at a recursion node with node-local coefficient-lattice exponent $\sigma$ (Definition 3.2), the normalized operands satisfy the dimension-free dot-product bounds $|T_{ij}|\leq 1$ and $|L_{ij}|\leq 1$.
@@
Corollary 7.2 (Integral mode: deterministic dyadic base with fixed slack and node certificate (no tuning knobs)).
Fix the slack constant $\delta_{0}:=2^{-3}=1/8$. In exact dyadic lattice semantics (Part I), it suffices to enforce the sufficient-base condition
$D^{2}2^{\sigma}\left(\frac{2}{\beta}+\frac{1}{\beta^{2}}\right)\;\leq\;\frac{1}{2}-\delta_{0},$ (55)
where $\sigma$ is the node-local coefficient-lattice exponent (Definition 3.2). For dyadic bases $\beta=2^{b}$ with $b\in\mathbb{Z}_{\geq 1}$, a fully deterministic choice with no tuning is: start with the candidate $b=t+3$ where $t:=\lfloor\log_{2}(D^{2}2^{\sigma})\rfloor$; if Eq. 55 fails, increment $b$ once (doubling $\beta$). This rule (and the corresponding exact integer certificate) is spelled out in Appendix  B.5.
Also, because $D^{2}2^{\sigma}$ is just a power-of-two shift of the fixed integer $D^{2}$, we have $t=\sigma+\lfloor\log_{2}(D^{2})\rfloor$, so the candidate $b=t+3$ is obtained by a single integer add once $\sigma$ is known.
Proof.
This is Theorem  7.1 specialized to the fixed slack $\delta=\delta_{0}$. The dyadic selection rule and the exact integer certificate are proved in Appendix  B.5. ∎
Remark 7.3 (Why a growing base is not “cheating” (and what $\sigma$ means)).
Condition (54) implies that $\beta$ must grow roughly in proportion to $2^{\sigma}$. There is no need for a “speed-first” policy: in exact (integral) dyadic mode we fix a small constant slack $\delta_{0}:=2^{-3}=1/8$ (not a tuning knob) and choose the dyadic base $\beta=2^{b}$ deterministically by Corollary 7.2 / Appendix  B.5, with a constant-cost integer certificate that the sufficient-base inequality holds at each node. Here $\sigma$ is the node-local coefficient-lattice exponent from Definition 3.2, i.e., it is determined by the smallest dyadic lattice step on which the normalized dot-product coefficients at that node actually lie. In worst-case analyses one may upper bound $\sigma$ by a quantity proportional to recursion depth (i.e., $O(\log n)$ in a $2\times 2$ recursion), but implementations MUST compute $\sigma$ from exponent-offset metadata of $A^{\prime},B^{\prime}$ rather than setting $\sigma:=$ (recursion depth). Consequently, the required bit-width of the dyadic base $\beta=2^{b}$ grows at most logarithmically in $n$ in the worst case. This is not a violation of the model; rather, it is the precise origin of the digit-band staging factor $K(n)$ (Theorem 9.5). The algorithm exchanges the uncontrolled magnitude growth of classical methods for a strictly controlled, logarithmic growth in precision (staging depth), giving the soft-quadratic complexity bound.
@@ 8.1 Centering correction identity
Let $A,B\in\mathbb{R}^{n\times n}$ and let $\alpha,\rho\in\mathbb{R}^{n}$, and define the centered matrices:
@@
$AB~=~A_{0}B_{0}+\mathbf{1}(\alpha^{\top}B_{0})+(A_{0}\rho)\mathbf{1}^{\top}+(\alpha^{\top}\rho)\mathbf{1}\mathbf{1}^{\top}.$ (58)
@@
From the definitions, $A=A_{0}+\mathbf{1}\alpha^{\top}$ and $B=B_{0}+\rho\mathbf{1}^{\top}$, so:
$AB$ $=$ $(A_{0}+\mathbf{1}\alpha^{\top})(B_{0}+\rho\mathbf{1}^{\top})$ (59)
$=$ $A_{0}B_{0}+A_{0}(\rho\mathbf{1}^{\top})+(\mathbf{1}\alpha^{\top})B_{0}+(\mathbf{1}\alpha^{\top})(\rho\mathbf{1}^{\top}).$ (60)
Now observe that:
$A_{0}(\rho\mathbf{1}^{\top})$ $=$ $(A_{0}\rho)\mathbf{1}^{\top},$ (61)
@@
$(\mathbf{1}\alpha^{\top})(\rho\mathbf{1}^{\top})$ $=$ $\mathbf{1}(\alpha^{\top}\rho)\mathbf{1}^{\top}$ (63)
$=$ $(\alpha^{\top}\rho)\mathbf{1}\mathbf{1}^{\top}.$ (64)
@@
We refer to this as the row(A)-col(B) policy. It implies the entrywise bound $|(A^{\prime}B^{\prime})_{ij}|\leq 1$ by Cauchy–Schwarz. This bound is the invariant used to justify the node-local dyadic base choice $\beta=\beta(\sigma)$ via Theorem 7.1 once the node-local lattice exponent $\sigma$ is computed from $A^{\prime},B^{\prime}$ (Definition 3.2).
@@
Centering enforces mean constraints. Normalization enforces norm constraints. These operations generally do not commute, and we cannot in general enforce zero-mean rows and columns and unit row and column norms simultaneously. Throughout, we use the row(A)–col(B) policy (Definition 8.2) to guarantee the entrywise bound $|(A^{\prime}B^{\prime})_{ij}|\leq 1$ needed for the node-local dyadic base choice $\beta=\beta(\sigma)$ once $\sigma$ is computed from exponent offsets (Definition 3.2).
@@
This choice is deliberate because it implies the entrywise bound $|(A^{\prime}B^{\prime})_{ij}|\leq 1$, which is the invariant needed to justify the node-local dyadic base choice $\beta=\beta(\sigma)$ once $\sigma$ is computed (Definition 3.2; Appendix B.5).
@@ 8.2 Implementation contract (integral kernel)
    No square roots: Dyadic $\ell_{2}$ normalization uses the lifted squared norm $S(x)\in\mathbb{Z}[1/2]_{\geq 0}$ and its integer lift $\widetilde{S}(x)\in\mathbb{Z}_{\geq 0}$ (Definition 6.1), and computes $k(x)$ via bitlength comparisons on $\widetilde{S}(x)$ (Proposition 6.3). Scaling by $2^{\pm k}$ is an exponent-offset update.
@@
    Sufficient base with computed $\sigma$ (no depth substitution). At each node, $\sigma$ MUST be computed from the exponent offsets of the normalized operands $A^{\prime},B^{\prime}$ as in Definition 3.2 (equivalently Eq. 266). Implementations MUST NOT set $\sigma$ equal to recursion depth (this can over-inflate $\beta$ and destroy performance). Given the computed $\sigma$, implementations MUST use the fixed slack convention $\delta_{0}:=1/8$ and set the dyadic packing base $\beta=2^{b}$ deterministically by Appendix B.5. A reference implementation MUST certify the sufficient-base inequality Eq. 263 for the chosen $\beta$. If it fails at the initial candidate, apply the one-step upgrade rule $b\leftarrow b+1$ (doubling $\beta$) and re-evaluate. If it still fails after the one-step upgrade, the implementation MUST take the failure-policy fallback of Appendix  B.5 (call ClassicalMM on the original subproblem and return) and MUST NOT apply the sieve at that node. (Equivalently, in staged-precision mode an implementation may signal failure to a retry wrapper and re-attempt with larger staging; see Algorithm  2.)
@@
    Shifts do not touch mantissas: Multiplication by $2^{\pm t}$ and by $\beta^{\pm 1}$ must be implemented as exponent updates, not as integer multiplication on mantissas.
@@ 8.3 Kernel pseudocode
(1) Center. Compute the exact column-mean vector $\alpha\in\mathcal{L}_{D}^{n}$ of $A$ and exact row-mean vector $\rho\in\mathcal{L}_{D}^{n}$ of $B$ using dyadic division:
@@
$\rho_{i}$ $:=$ $\frac{1}{n}\sum_{j=1}^{n}B_{ij}.$
Set $A_{0}:=A-\mathbf{1}\alpha^{\top}$ and $B_{0}:=B-\rho\mathbf{1}^{\top}$.
@@
(2b) Compute the node-local lattice exponent $\sigma$ (no extra passes). Using exponent offsets $\exp(\cdot)$ in the representation $x=(w/D)2^{e}$ (Appendix B.1), compute:
$a_{r}$ $:=$ $\min_{i}\exp(A^{\prime}_{ir}),~b_{r}:=\min_{j}\exp(B^{\prime}_{rj}),$ (69)
$\mathcal{R}$ $:=$ $\{\,r:a_{r}<\infty~\text{and}~b_{r}<\infty\,\}.$ (70)
If $\mathcal{R}=\emptyset$, set $\sigma:=0$; otherwise set $g_{\min}:=\min_{r\in\mathcal{R}}(a_{r}+b_{r})$ and $\sigma:=\max\{0,-g_{\min}\}$. This is exactly Definition 3.2 / Eq. 266, and can be accumulated during Step (2).
(3) Pack, recurse, and extract quadrants. Partition $A^{\prime},B^{\prime}$ into $2\times 2$ blocks of size $(n/2)\times(n/2)$.
/* Certified base selection (no tuning knobs; at most one upgrade). */
Compute the candidate exponent $b_{0}$ (i.e., $\beta_{0}=2^{b_{0}}$) from Eq. 270. Set $b\leftarrow b_{0}$ and $\beta\leftarrow 2^{b}$
Evaluate the exact integer certificate Eq. 271 (and, in a reference implementation, assert the normalization bounds from Step (2))
if the certificate fails at $b=b_{0}$ then
 $b\leftarrow b_{0}+1$$\beta\leftarrow 2^{b}$
 // one-step upgrade, Eq. 272
   Re-pack using the upgraded $\beta$ and re-evaluate the certificate (in exact integer arithmetic this second check must succeed; if it does not, treat it as overflow/truncation)
 if the certificate still fails then
    return ClassicalMM(A,B)(A,B)
    // exact fallback; Appendix B.5
foreach $(i,j)\in\{1,2\}^{2}$ do
@@
$C~:=~A_{0}B_{0}+\mathbf{1}(\alpha^{\top}B_{0})+(A_{0}\rho)\mathbf{1}^{\top}+(\alpha^{\top}\rho)\mathbf{1}\mathbf{1}^{\top}.$